### v1

# Inverse thermodynamic uncertainty relation and entropy production

Abstract
Nonequilibrium current fluctuations represent one of the central topics in nonequilibrium physics. The thermodynamic uncertainty relation (TUR) is widely acclaimed for rigorously establishing a lower bound on current fluctuations, expressed in terms of the entropy production rate and the average current. In this study, we focus on an upper bound for the fluctuations, referred to as the inverse thermodynamic uncertainty relation (iTUR). We derive a universal iTUR expression in terms of the entropy production rate for continuous-variable systems governed by over-damped Langevin equations, as well as for discrete-variable systems described by Markov jump processes. The iTUR establishes a no-go theorem prohibiting perpetual superdiffusion in systems with a finite entropy production rate and a finite spectral gap. The divergence of the variance of any current becomes possible only when the spectral gap vanishes or the entropy production rate diverges. As a relevant experimental scenario, we apply the iTUR to the phenomenon of giant diffusion, emphasizing the pivotal roles of the spectral gap and entropy production.
Introduction.— Understanding nonequilibrium fluctuations is a cornerstone of nonequilibrium statistical physics. Among the numerous breakthroughs in the field of stochastic thermodynamics [1, 2], the thermodynamic uncertainty relation (TUR) stands out as a landmark result, establishing a fundamental constraint on nonequilibrium current fluctuations and entropy production [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. The TUR is valid in the overdamped dynamics, and provides critical insights and applications in nonequilibrium thermodynamics [34, 35, 36, 37], such as a trade-off between precision and thermodynamic cost, and a measurable lower bound for entropy production, inspiring studies aimed at inferring entropy production from observable quantities [38, 39, 40, 41, 42, 43]. In particular, when interpreted as a lower bound on nonequilibrium fluctuations, the TUR takes the following form:
$D_{\infty}\ \geq\frac{j_{\rm st}^{2}}{\sigma_{\text{st}}}\,,$ (1)
where $D_{\infty}$ is the long-time fluctuations of the current defined as in the same way as the diffusion coefficient (detailed below in Eq. (6)), $j_{\rm st}$ is its steady state average, and $\sigma_{\text{st}}$ is the entropy production rate with the Boltzmann constant set to unity.
Beyond the TUR’s lower bound on nonequilibrium fluctuations, a significant advancement has recently been achieved in identifying the upper bound on the fluctuations. In Ref. [44], concentration inequality techniques in continuous-time Markov jump processes was employed to reveal that nonequilibrium fluctuations are bounded above by the instantaneous fluctuation and the spectral gap of the symmetrized transition rate matrix. This complementary result, also known as the inverse TUR (iTUR), has been extended to the quantum regime for counting observables and their first passage times [45, 46]. However, the derived upper bounds do not account for the role of thermodynamic cost such as entropy production and are restricted to discrete-variable. These limitations give rise to three pressing questions: Can we upper-bound the fluctuations in terms of entropy production to elucidate the role of thermodynamic costs? Is such an iTUR applicable to continuous-variable systems governed by overdamped Langevin dynamics? Can we use iTUR to obtain new insights into the mechanisms of physical phenomena?
In this study, we address these questions. Employing the variational principle for the cumulant generating function of the current, we derive a concise expression of iTUR that incorporates the entropy production rate for continuous variables (Eqs. (9) and (10) below), and for discrete variables (Eq. (14)). These iTURs elucidate the interplay between the entropy production rate and the spectral gap in governing current fluctuations. Large current fluctuations necessitate either a high entropy production rate or a small spectral gap. In particular, the iTUR establishes a no-go theorem on perpetual superdiffusion in systems with a finite entropy production rate and a non-zero spectral gap. Furthermore, the relations show that in equilibrium states, instantaneous fluctuations ($D_{0}$ in Eq. (6)) always exceed the long-time fluctuation ($D_{\infty}$).
We apply the iTUR to the phenomenon of giant diffusion observed in particles subjected to nonconservative forces within a periodic potential [47, 48, 49]. This phenomenon arises from the competition between local diffusion near potential minima and transitions across potential barriers driven by nonconservative forces, leading to a significant enhancement of the long-time diffusion constant at a critical nonconservative force. Experimentally, giant diffusion has been observed in systems such as biological rotary motors, such as F1-ATPase [50], colloidal particles [51] and DNA molecules [52, 53]. The iTUR provides a perspective on this phenomenon through the interplay of entropy production rate and spectral gap. Driving the system further out of equilibrium through an increased entropy production rate consistently enhances diffusion. This enhancement becomes more pronounced when the spectral gap is small, allowing fluctuations to persist for longer time. Our results illuminate how nonequilibrium driving enhances diffusion within the constraints of entropy production and relaxation properties.
The iTURs presented in this paper offer profound insights into the physical principles underlying nonequilibrium current fluctuations, framed through the lens of thermodynamic costs, particularly entropy production. Furthermore, it holds significant potential for applications that span a broad spectrum of nonequilibrium physical phenomena.
Setup for continuous variables.— We consider particles undergoing overdamped Langevin dynamics in $\mathbb{R}^{n}$ with a uniform temperature, described by the following stochastic differential equation
$\dot{\bm{x}}_{t}=\bm{A}(\bm{x}_{t})+\sqrt{2\bm{B}}\,\bm{\xi}(t)\,,$ (2)
where $\dot{\bm{x}}_{t}$ denotes the velocity vector of the particle, $\bm{A}(\bm{x})$ is the drift vector field, $\bm{B}$ is a positive-definite diffusion matrix, and $\bm{\xi}(t)$ represents Gaussian white noise with zero mean and correlations $\langle\xi_{i}(t)\xi_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$. The corresponding probability density function $p(\bm{x},t)$ evolves in time according to the Fokker–Planck equation [54, 55]:
$\begin{split}\partial_{t}p(\bm{x},t)&=\mathbb{L}p(\bm{x},t)\,,\\ \mathbb{L}p(\bm{x},t)&=-\nabla\cdot\left[\bm{A}(\bm{x})p(\bm{x},t)-\bm{B}\nabla p(\bm{x},t)\right].\end{split}$ (3)
In steady state, $\partial_{t}p_{\text{st}}(\bm{x})=0$, and the system may sustain persistent currents due to non-equilibrium conditions.
We define and consider the accumulation of the general current $\hat{\cal J}$ linearly dependent on the velocity with the weight function $\bm{\theta}(\bm{x})$ as
[…]
where $\circ$ denotes the Stratonovich product. The steady-state average current $j_{\rm st}$ is computed as $j_{\rm st}=\langle\hat{\cal J}\rangle/\tau$ within the framework of the Langevin dynamics. In the Fokker-Planck picture, an equivalent expression is
[…]
where $\bm{\nu}_{\text{st}}({\bm{x}}):=\bm{A}({\bm{x}})-\bm{B}\nabla\ln p_{\text{st}}({\bm{x}})$ is local mean velocity. Note that the product ${\bm{\nu}}_{\text{st}}({\bm{x}})p_{\text{st}}({\bm{x}})$ is the probability current at the steady state. We also introduce the fluctuation of the current over a time interval $\tau$, defined as
$D_{\tau}=\frac{\text{Var}(\hat{\cal J})}{2\tau}\,,$ (6)
Here, ${\rm Var}(\hat{\cal J})$ represents the variance of the accumulated stochastic current given by Eq. (4). Inclusion of the factor $2$ in the denominator ensures consistency with the standard definition of the diffusion coefficient, particularly when the weight function ${\bm{\theta}}$ is set to all-ones vector.
As a fundamental measure of thermodynamic cost, we consider the entropy production rate $\sigma_{\text{st}}$ at the steady state, expressed as [56]:
[…]
This expression explicitly demonstrates the nonnegativity of the entropy production rate. For steady-state thermal systems, the entropy production rate expresses the rate of heat dissipation. Throughout this paper, we set the Boltzmann constant to unity.
The TUR establishes a lower bound on the fluctuations $D_{\tau}$ in terms of the entropy production rate and the steady-state current. A more general result than Eq. (1) is the finite-time TUR, which provides a bound on the fluctuations in the steady state at arbitrary time interval [57, 15]:
$D_{\tau}\geq\frac{j_{\text{st}}^{2}}{\sigma_{\text{st}}}=:D_{\tau,{\bm{\theta}}}^{\textrm{TUR}}\,.$ (8)
To establish an upper bound on fluctuations, the iTUR was formulated using the spectral gap of a symmetrized generator of the original Markov jump process [44]. In this work, we extend this concept to the continuous-variable setting by considering the spectral gap $\lambda$ of the symmetrized Fokker–Planck operator, which is given by replacing the drift term ${\bm{A}}({\bm{x}})$ by $\tilde{\bm{A}}(\bm{x})=\bm{B}\nabla\ln p_{\text{st}}(\bm{x})$ in the operator $\mathbb{L}$. Importantly, the dynamics governed by this operator preserves the same steady state $p_{\text{st}}(\bm{x})$. The spectral gap is related to the slowest timescale of the equilibrium relaxation with the same diffusion matrix $\bm{B}$ and the steady state $p_{\text{st}}(\bm{x})$. While symmetrization alters the dynamics, the spectral gap remains closely linked to the relaxation properties of the original dynamics (see, for example, [58]).
Main results for the continuous variable case.— We present here the main results for continuous variables governed by the dynamics (3). An outline of the derivation is provided at the end of this paper, with details presented in the supplementary material (SM)[59]. First, we present the result for the most typical and experimentally relevant scenario, namely, the standard diffusion case with ${\bm{\theta}}$ set to the all-ones vector. We derive the bound on the long-time diffusion coefficient as
$D_{\infty}\leq D_{0}+\frac{D_{0}\sigma_{\text{st}}-j_{\text{st}}^{2}}{\lambda}=:D_{\infty}^{\rm iTUR}\,,~{}~{}(\theta({\bm{x}})=1\,,\forall{\bm{x}})\,,$ (9)
where the non-negativity of $D_{0}\sigma_{\text{st}}-j_{\text{st}}^{2}$ follows directly from the short-time limit of the TUR (8). The iTUR yields $D_{\infty}/D_{0}\leq 1+\sigma_{\text{st}}/\lambda$, elucidating how the thermodynamic cost $\sigma_{\rm st}$ and the relaxation rate through $\lambda$ jointly constrain the long-time diffusion coefficient. A large long-time diffusion constant can only arise either through a high entropy production rate or a small spectral gap. A smaller $\lambda$ physically allows correlations to persist for longer durations, thereby increasing the upper bound on the long-time diffusion constant. In the limit of very large $\lambda$, the bound approaches $D_{0}$, demonstrating that rapid relaxation negates enhanced diffusion.
Second, we derive the result for finite measurement times $\tau$ at the steady state for a general current, i.e., a current with a general weight function:
$D_{\tau}\leq D_{0}+g(\bar{\lambda}\tau)\frac{D_{\text{max}}\sigma_{\text{st}}-j_{\text{st}}^{2}}{\lambda}=:D_{\tau,{\bm{\theta}}}^{\rm iTUR},~{}~{}~{}~{}~{}({\bm{\theta}}:{\rm general}).$ (10)
The detailed proof is given in the SM [59]. Here, $D_{\text{max}}$ and $\sigma_{\text{max}}$ represent the maximum value of the instantaneous current fluctuation and entropy production, satisfying $\bm{\theta}\cdot\bm{B}\bm{\theta}\leq D_{\text{max}}$ and $\bm{\nu}_{\mathrm{st}}\cdot\mathbf{B}^{-1}\bm{\nu}_{\mathrm{st}}\leq\sigma_{\text{max}}$, respectively. The characteristic rate $\bar{\lambda}:=\sqrt{\lambda(\lambda+\sigma_{\text{max}})}$ sets a timescale. The function $g(x)$ is defined as
[…]
which is nonnegative, monotonically increasing, and smoothly interpolates between 00 and $1$ as $x$ changes from 00 to +$\infty$+ ∞. In the limit of the vanishing measurement time $\tau\to 0$ with $\sigma_{\text{max}}$ finite, the bound simplifies to $D_{\tau}\leq D_{0}$, which is achieved with equality. In contrast, as $\tau\to\infty$, the function $g(\bar{\lambda}\tau)$ approaches $1$, and hence the inequality (10) covers the specific case (9) by setting the weight function to all-ones vector.
The iTUR for the general case (10) elucidates similar physics to the specific case (9) regarding the roles of the entropy production rate and spectral gap for large current fluctuations. Moreover, the rigorous relation $D_{\tau}\leq D_{0}$ is established in the equilibrium states for any measurement time. The TUR and iTUR constrain the range of current fluctuations for arbitrary measurement time as
$D_{\tau,{\bm{\theta}}}^{\rm TUR}$ $\leq D_{\tau}\leq D_{\tau,{\bm{\theta}}}^{\rm iTUR}\,.$ (12)
When considering particle diffusion setting $\theta({\bm{x}})=1$, $D_{\tau}$ behaves as $D_{\tau}\sim\bar{D}\tau^{\nu}$ at large scale $\tau$, where $\nu<0\,(>0)$ implies sub(super) diffusion. The TUR imposes a constraint on the cutoff time marking the end of subdiffusion [60], while the iTUR similarly limits the duration of superdiffusive behavior. Notably, the iTUR establishes a no-go theorem, ruling out the possibility of perpetual superdiffusion in systems with a finite entropy production rate and a finite spectral gap.
Application to the giant diffusion.— We apply the iTUR for the diffusion coefficient (9) to the giant diffusion phenomenon, which is characterized by the emergence of anomalously large diffusion coefficients exceeding the local diffusion constant in systems governed by periodic potentials under the influence of nonconservative forces [47, 48]. Giant diffusion has been experimentally observed in various biological and soft matter systems [50, 51, 52, 53]. The iTUR (9) establishes that such enhanced diffusivity is achievable only under conditions of either a substantial entropy production rate or a small spectral gap. Within this theoretical framework, we consider this phenomenon. The dynamics of a particle in this system is governed by the stochastic differential equation
$\dot{x}_{t}=-V^{\prime}(x_{t})+F+\sqrt{2D_{0}}\,\xi(t)\,,$ (13)
where $V(x)=V_{0}\sin\bigl{[}(2\pi x)/L\bigr{]}$ is a periodic potential with amplitude $V_{0}$ and period $L$, $F$ is a constant nonconservative force and $D_{0}=T/\gamma$ is the local diffusion coefficient determined by the temperature $T$ and friction coefficient $\gamma$. In our analysis, we set the friction coefficient $\gamma$ to unity. To discuss the long-time diffusion constant, we define the normalized diffusion constants:
$\begin{split}\tilde{D}_{\infty}&:=D_{\infty}D_{0}^{-1}\,,\\ \tilde{D}_{\infty}^{{\rm iTUR}/{\rm TUR}}&:=D_{\infty}^{{\rm iTUR}/{\rm TUR}}D_{0}^{-1}\,,\end{split}$
where $D_{\infty}^{\rm TUR}=D_{\infty,{\bm{\theta}}}^{\rm TUR}$ with $\theta({\bm{x}})=1,\forall{\bm{x}}$.
Figure 1: Diffusion coefficient, iTUR and TUR as functions of the external force $F$ in a tilted periodic potential. Parameters are $V_{0}=1$, $L=1$, and $D_{0}=0.5(T=0.5)$. The dashed line represents $\tilde{D}_{\infty}$, showing a significant increase near the critical force $F_{c}=2\pi$, indicative of giant diffusion. The solid line denotes the iTUR upper bound, closely aligning with $\tilde{D}_{\infty}$, while the dotted line represents the TUR lower bound. The inset shows the behavior of the entropy production rate $\sigma_{\text{st}}$ and spectral gap $\lambda$ with $F$. The left y-axis corresponds to the spectral gap, while the right y-axis represents the entropy production rate.
Fig. 1 shows the dependence of the normalized diffusion coefficients on the external force $F$. As $F$ increases, $\tilde{D}_{\infty}$ exhibits a pronounced enhancement near the critical force $F_{c}\sim(2\pi V_{0})/L$, corresponding to the point where the potential barrier vanishes in the infinite-line picture. The iTUR provides an upper bound for $\tilde{D}_{\infty}$, as evidenced by the close agreement between the theoretical prediction of the iTUR (solid line) and the observed diffusion coefficients (dashed line). In contrast, the TUR establishes a complementary lower bound based on entropy production rate. The inset illustrates the $F$-dependence of the spectral gap $\lambda$ and the entropy production rate $\sigma_{\text{st}}$. As $F$ approaches $F_{c}$, the spectral gap decreases, indicating slower relaxation dynamics, while the entropy production rate begins to increase near this point. Slower relaxation (smaller $\lambda$) allows fluctuations to persist for longer durations, whereas higher entropy production rate drives enhanced transport. This interplay aligns with the behavior predicted by the upper bound by the iTUR. For this specific potential, the iTUR accurately captures diffusion enhancement up to the critical force $F_{c}$ but does not explain the subsequent decrease beyond it, whereas in a generic tilted potential, the iTUR might predict a peak at $F_{c}$.
Figure 2: Dependence of the diffusion coefficient $D_{\infty}$ and the iTUR upper bound on the entropy production rate $\sigma_{\text{st}}$ for fixed spectral gaps $(\lambda=5,10,100)$. In both panels, dashed lines represent $D_{\infty}/D_{0}$, and solid lines denote the iTUR upper bound.
Fig. 2 shows the normalized diffusion constant as a function of entropy prodction rate $\sigma_{\text{st}}$ for a fixed spectral gap $\lambda$. We observe that $\tilde{D}_{\infty}$ increases monotonically with an increasing entropy production rate. The figure demonstrate that pushing the system further out of equilibrium via increased entropy production rate consistently amplifies diffusion, and this effect is more pronounced at smaller spectral gaps.
Discrete variable case.— We consider a continuous-time Markov jump process that describes transitions between discrete states. The transition rates $W_{km}$ from state $m$ to state $k$ define these transitions. To ensure thermodynamic consistency, we assume that the transition rates satisfy the local detailed balance condition. In the steady state, the state probabilities $\pi_{k}$ remain constant over time. Let $\omega_{\tau}=\{n_{0},(n_{1},t_{1}),\ldots,(n_{N},t_{N})\}$ be a stochastic trajectory of the system during the time interval $[0,\tau]$, where the system is initially in state $n_{0}$ and a transition from state $n_{i-1}$ to state $n_{i}$ occurs at time $t_{i}$ for each $1\leq i\leq N$. For each trajectory, we consider a arbitrary current observable $\hat{\cal J}(\omega_{\tau})=\sum_{i=1}^{N}d_{n_{i}n_{i-1}}$, where the increment $d_{mn}$ associated with the transition $n\to m$, satisfying $d_{mn}=-d_{nm}$. The steady-state time average of this observable is then given by $j_{\text{st}}=\sum_{k\neq m}d_{km}W_{km}\pi_{m}$. The entropy production rate is expressed as: $\sigma_{\text{st}}=\sum_{k>m}J_{km}\ln\bigl{[}(W_{km}\pi_{m})/(W_{mk}\pi_{k})\bigr{]}$, where $J_{km}=W_{km}\pi_{m}-W_{mk}\pi_{k}$ are the steady-state fluxes. Let $\kappa=\max_{k}\sum_{m\neq k}W_{km}$, which represents the maximum escape rate; and the spectral gap $\lambda$ of the symmetrized matrix $(W^{\top}+W^{\dagger})/2$ where $W^{\dagger}$ is the adjoint of $W^{\top}$ taken with respect to the inner product induced by the stationary stationary state $\pi$. For a generalized current with $d_{km}=-d_{mk}\leq 1$, we derive the iTUR for the Markov jump process:
$D_{\tau}\leq D_{0}+g(\tilde{\lambda}\tau)\frac{\kappa\sigma_{\text{st}}-2j_{\text{st}}^{2}}{2\lambda}\,,$ (14)
where $\tilde{\lambda}:=\sqrt{\lambda(\lambda+2\kappa)}$. Note that the iTURs for the discrete variable case cannot reproduce the iTURs for continuous variable case since the escape rate $\kappa$ diverges in the continuous variable limit. In the SM [59], we provide the proof for a tighter iTUR [61].
Outline of the proof.— We here outline the proof, while the details are presented in the SM [59]. Utilizing a variational principle for diffusion coefficients, we can express $D_{\infty}$ as
$D_{\infty}$ $=\sup_{\bm{\chi}}\inf_{\bm{\eta}}\bigg{[}2\,\text{Cov}_{\text{st}}\left((\bm{\theta}+\nabla\bm{\eta})\cdot\bm{\nu}_{\text{st}},\bm{\chi}\right)$ (15)
$\quad+\langle(\bm{\theta}+\nabla\bm{\eta})\cdot\bm{B}(\bm{\theta}+\nabla\bm{\eta})\rangle_{\text{st}}-\langle\nabla\bm{\chi}\cdot\bm{B}\nabla\bm{\chi}\rangle_{\text{st}}\bigg{]},$
where $\langle\diamond\rangle_{\text{st}}$ and $\text{Cov}_{\text{st}}(\diamond,\diamond)$ denote the steady-state average and covariance, respectively. Setting $\bm{\eta}(\bm{x})=\bm{0}$ simplifies Eq. (15) to
$D_{\infty}$ $\leq D_{0}+\sup_{\bm{\chi}}\left[2\,\text{Cov}_{\text{st}}(\bm{\theta}\cdot\bm{\nu}_{\text{st}},\bm{\chi})-\langle\nabla\bm{\chi}\cdot\bm{B}\nabla\bm{\chi}\rangle_{\text{st}}\right],$ (16)
where $D_{0}=\langle\bm{\theta}\cdot\bm{B}\bm{\theta}\rangle_{\text{st}}$. Using the Cauchy–Schwarz inequality, we obtain
$D_{\infty}\leq D_{0}+\sup_{\bm{\chi}}\left[\frac{\left[\text{Cov}_{\text{st}}(\bm{\theta}\cdot\bm{\nu}_{\text{st}},\bm{\chi})\right]^{2}}{\langle\nabla\bm{\chi}\cdot\bm{B}\nabla\bm{\chi}\rangle_{\text{st}}}\right].$ (17)
We define the symmetrized Fokker–Planck operator, $\mathbb{L}_{\rm sym}$, by substituting the original drift in $\mathbb{L}$ with the modified field $\tilde{\bm{A}}(\bm{x})=\bm{B}\nabla\ln p_{\text{st}}(\bm{x})$. This substitution renders $\mathbb{L}_{\rm sym}$ symmetric with respect to a particular inner product weighted by the steady-state distribution. Its spectral gap, $\lambda$, can then be determined via the variational formula:
$\lambda=\inf_{\chi}\left[\frac{\langle\nabla\chi\cdot\bm{B}\nabla\chi\rangle_{\text{\rm st}}}{\text{Var}_{\text{\rm st}}(\chi)}\right].$ (18)
We can deduce that $\lambda$ is upper bounded by the ratio $\langle\nabla\chi\cdot\bm{B}\nabla\chi\rangle_{\text{st}}/\text{Var}_{\text{st}}(\chi)$. This leads to the bound:
$D_{\infty}\leq D_{0}+\frac{1}{\lambda}\text{Var}_{\text{st}}(\bm{\theta}\cdot\bm{\nu}_{\text{st}})\,.$ (19)
We now focus on bounding $\text{Var}_{\text{st}}(\bm{\theta}\cdot\bm{\nu}_{\text{st}})$. By definition, we have
$\text{Var}_{\text{\rm st}}(\bm{\theta}\cdot\bm{\nu}_{\text{\rm st}})=\langle(\bm{\theta}\cdot\bm{\nu}_{\text{\rm st}})^{2}\rangle_{\text{\rm st}}-j^{2}_{\text{\rm st}}\,.$ (20)
The first term can be bounded as follows:
$\langle(\bm{\theta}\cdot\bm{\nu}_{\text{\rm st}})^{2}\rangle_{\text{\rm st}}\leq\langle(\bm{\theta}\cdot\bm{B}\bm{\theta})(\bm{\nu}_{\text{\rm st}}\cdot\bm{B}^{-1}\bm{\nu}_{\text{\rm st}})\rangle_{\text{\rm st}}\,.$ (21)
Introducing $D_{\text{max}}$, satisfying $\bm{\theta}\cdot\bm{B}\bm{\theta}\leq D_{\text{max}}$, and noting $\sigma_{\text{st}}=\langle\bm{\nu}_{\text{st}}\cdot\bm{B}^{-1}\bm{\nu}_{\text{st}}\rangle_{\text{st}}$, we obtain
$\text{Var}_{\text{st}}(\bm{\theta}\cdot\bm{\nu}_{\text{st}})\leq D_{\text{max}}\sigma_{\text{st}}-j_{\text{st}}^{2}\,.$ (22)
Substituting this bound into Eq. (19), we arrive at the iTUR bound (10) with $\tau\to\infty$.
Summary and Discussion.— We have derived upper bounds on current fluctuations at steady state (iTUR) by utilizing both the entropy production rate and the spectral gap. By integrating the thermodynamic uncertainty relation (TUR), our results constrain the range of current fluctuations. While the TUR suggests that current fluctuations relative to the average current can be reduced by increasing the entropy production rate, our iTUR involving entropy production implies that the overall extent of fluctuations tends to increase. Crucially, this effect is not apparent in the iTUR derived in Ref. [44]. We apply the derived iTUR to the phenomenon of giant diffusion, demonstrating the critical roles of both the spectral gap and the entropy production rate.
The iTURs introduced in this paper have significant potential for broad applications across relevant systems. For instance, the iTUR provides thermodynamic concentration inequalities, imposing upper limits on the probability of observing current deviations significantly far from their mean. Applications include not only enhancing our understanding of the underlying mechanisms, but also facilitating the estimation of maximum fluctuations and the design of efficient transport performance.
Acknowledgments
This work is supported by JSPS KAKENHI Grant No. JP23K25796. AD is supported by JSPS KAKENHI (Grant No. 22K13974 and 24H00833).